OpenStudy (anonymous):

A square garden plot measures 125 square feet. A second square garden plot measures 405 square feet. How many more feet of fence will the second garden plot require over that required for the first garden plot? A) 16 (square root)5 feet of additional fencing B) 16 (square root)3 feet of additional fencing C) 80 feet of additional fencing D) 48 feet of additional fencing

OpenStudy (phi):

A square garden plot measures 125 square feet. can you find the length of the side of this square ?

OpenStudy (anonymous):

I wish that's all the word problem has for my homework lol..... like I said I hate word problems no matter how I try solving it I don't get any of those answers... so annoying

OpenStudy (phi):

A square garden plot measures 125 square feet. can you find the length of the side of this square ?

OpenStudy (anonymous):

side of first garden =sqrt(125)=11.18 ft Fencing required =4s=4*11.15 =44.72ft side of second garden =sqrt(405)=20.12 ft Fencing required =4s=4*20.12=80.48 ft Extra fencing required for second garden =80.48-20.12 =60.36 ft. but that's not one of the answers.... what am I doing wrong

OpenStudy (anonymous):

\[\sqrt{405}*4-\sqrt{125}*4=16 \sqrt{5} \]

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

wow... I feel dumb.. I was so over working it..... YOU ROCK

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Thank you for helping me score a 100!!!! I been stuck on that question for 45mins